Write the equation of the line that passes through the points (6, 2) and (6, 10).

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Write the equation of the line that passes through the points (6, 2) and (6, 10).
----------
x is 6 in both points.
--> x = 6

Answer by stanbon(75887)

Write the equation of the line that passes through
the points (6, 2) and (6, 10).
-----
slope = (10-2)/(6-6) = 8/0 is undefined
------
Since the x-values are the same the equation is x = 6
=======================
Plot the points and draw a line thru them so you can see the answer.
======================================================================
Cheers,
Stan H.